OpenStudy (amber15):

The force needed to keep a car from skidding on a curve varies inversely as the radius of the curve and jointly as the weight of the car and the square of the speed . if 210 lb of force keeps a 2,500 lb car from skidding on a curve of radius 500 ft at 25 mph . what force would keep the same car from skidding on a curve of radius 750 ft at 40 mph

OpenStudy (amber15):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what is the initial equation? were you able to figure that out?

OpenStudy (amber15):

no i'm lost on this

OpenStudy (amber15):

i really forgot how to do this one

jimthompson5910 (jim_thompson5910):

ok we are going to have these variables F = force needed to keep a car from skidding on a curve r = radius w = weight of the car s = speed of the car

OpenStudy (amber15):

ok

jimthompson5910 (jim_thompson5910):

"The force needed to keep a car from skidding on a curve varies inversely as the radius of the curve" that allows us to say F = k/r

jimthompson5910 (jim_thompson5910):

tack on "jointly as the weight of the car and the square of the speed" and you'll have \[\Large F = \frac{k*w*s^2}{r}\] where k is some constant

OpenStudy (amber15):

ok

jimthompson5910 (jim_thompson5910):

now we get to the part " 210 lb of force keeps a 2,500 lb car from skidding on a curve of radius 500 ft at 25 mph" F = 210 w = 2500 r = 500 s = 25 plug those values into \[\Large F = \frac{k*w*s^2}{r}\] and solve for k

OpenStudy (amber15):

ok

OpenStudy (amber15):

k=42/625

OpenStudy (amber15):

its looking for lb

jimthompson5910 (jim_thompson5910):

k = 42/625 = 0.0672 good

jimthompson5910 (jim_thompson5910):

"what force would keep the same car from skidding on a curve of radius 750 ft at 40 mph" so it wants the value of F where k = 0.0672 w = 2500 r = 750 s = 40

OpenStudy (amber15):

210

jimthompson5910 (jim_thompson5910):

that was the old value of F

OpenStudy (amber15):

750

jimthompson5910 (jim_thompson5910):

plug k = 0.0672 w = 2500 r = 750 s = 40 into \[\Large F = \frac{k*w*s^2}{r}\]

OpenStudy (amber15):

268.800

OpenStudy (amber15):

oops 537.6.

jimthompson5910 (jim_thompson5910):

\[\Large F = \frac{k*w*s^2}{r}\] \[\Large F = \frac{0.0672*2500*40^2}{750}\] \[\Large F = \frac{268,800}{750}\] \[\Large F = 358.4\]

OpenStudy (amber15):

oh okay

OpenStudy (amber15):

i finally caught on how this was done Thank you